OpenStudy (anonymous):

a rock is thrown upward from the edge of a 50 m cliff overlooking a lake, with an initial velocity of 17.2 m/s. Define variables and write and equation that models the height of the rock.

OpenStudy (anonymous):

we will consider only the vertical direction use the equation: \[y_2(t) - y_1 = V_0 t + \frac{ 1 }{ 2 }a t^2\] where y2 and y1 are the heights at a time 2 and a time 1 V_0 is the initial veloity a is acceleration [gravity in our case = -9.8 m/s^2] t is the difference in time between height 1 and 2

OpenStudy (anonymous):

in our model, y_1 is the initial height [50], so simplified, our model is: \[y_2(t)=50+17.2t−4.4t^2\]

OpenStudy (anonymous):

where does the 4.4t^2 come from?

OpenStudy (anonymous):

wow, made a mistake. should be -9.8/2 = 4.9

OpenStudy (anonymous):

\[y_2(t)=50+17.2t−4.9t^2\]

OpenStudy (anonymous):

where does the -9.8 come from?

OpenStudy (anonymous):

gravity = -9.8 m/s^2 = acceleration on rock

OpenStudy (anonymous):

oh ok thanks!